find two consecutive odd integers such that five times the first integer is 25 less than twice the greater integer
Found 2 solutions by stanbon, ewatrrr:
Answer by stanbon(75887)

find two consecutive odd integers such that five times the first integer is 25 less than twice the greater integer
----
1st: 2x-1
2nd: 2x+1
---
Equation:
5(2x-1) = 2(2x+1)-25
10x-5 = 4x+2-25
6x = -18
x = -3
-----
1st: 2x-1 = -7
2nd: 2x+1 = -5
=========================
Cheers,
Stan H.
==============

Answer by ewatrrr(24785)


Hi,
Let x represent the smaller of the two consecutive odd integers.
the other then would be (x+2)
Write as we Read
5x = 2(x+2) - 25
solving for x
5x = 2x + 4 - 25
3x = -21
x = -7, the smaller consecutive odd integers. The larger is -5 (-7+2)
Checking our answer
5*-7 = 2*-5 - 25 = 35